Demelza purchased her favorite perfume three months ago, J’adore by Dior, for $75. She logs online to purchase one for her friends birthday and it is now $100. What is the percentage change in the perfume price?

Select one:

a. 75%

b. 50%

c. 25%

d. 33%
Answer #1

Initial Price=Po=$75

Final Price=P1=$100

Percent change in perfume price=(P1-Po)/Po=(100-75)/75=33.33%

Correct option is

d. 33%